### v1

Abstract
Deploying a Machine Learning (ML) training pipeline into production requires robust software engineering practices. This differs significantly from experimental workflows. This experience report investigates this challenge in SPIRA, a project whose goal is to create an ML-Enabled System (MLES) to pre-diagnose insufficiency respiratory via speech analysis. The first version of SPIRA’s training pipeline lacked critical software quality attributes. This paper presents an overview of the MLES, then compares three versions of the architecture of the Continuous Training subsystem, which evolved from a Big Ball of Mud, to a Modular Monolith, towards Microservices. By adopting different design principles and patterns to enhance its maintainability, robustness, and extensibility. In this way, the paper seeks to offer insights for both ML Engineers tasked to productionize ML training pipelines and Data Scientists seeking to adopt MLOps practices.
Keywords:
Code Quality MLOps Software Architecture Machine Learning Enabled Systems Healthcare Domain Case Study
1 Introduction
In 2020, amidst the COVID-19 pandemic, a multidisciplinary team of researchers from the University of São Paulo (USP) created SPIRA  [^1] https://github.com/spirabr: a project to detect respiratory insufficiency via speech analysis, using Machine Learning (ML) [5].
Since then, the scope of SPIRA has evolved: its proposal is to detect respiratory insufficiency of different origins. Multiple sickness can cause this symptom, including smoking side effects, flu, severe asthma, and heart conditions [5].
To create a tool that could assist physicians, the team proposed to develop the SPIRA ML-Enabled System (MLES). Since 2020, different components have been incrementally developed by bachelor students working with the project.
The paper is divided as follows. Section 2 presents the architecture of the MLES. Section 3 refers to previous research and narrows the scope to incremental development of the Continuous Training subsystem. Sections 4 to 6 explain how its architecture evolved from a Big Ball of Mud [6] to a Modular Monolith [15] to Microservices [14]. Sections 7 and 8 lists challenges and lessons learned. Finally, section 9 summarizes contributions and future work.
2 The SPIRA ML-Enabled System
This section describes the architecture of the SPIRA ML-Enabled System, as illustrated by figure 1. Sections 2.1 to 2.6 showcase its six subsystems according to the reference architecture extended by Ferreira et al. [3].
Figure 1: Architecture of the SPIRA ML-Enabled System. This architecture follows the same notation presented in Ferreira et al. [3]. Rectangles represent services, which execute continuously. Stacked rectangles represent pipelines, which execute a task on demand. Lastly, cylinders represent data storage, which may be databases of any type. Components are connected by arrows. Black arrows with a hollow tip illustrate the execution flow. They start and end in a component. Labeled arrows represent the trigger that starts a workflow, whereas unlabeled arrows represent the continuation of an existing workflow. Colored arrows with a filled tip illustrate the data flow. They appear in two types: solid arrows going to and from a data storage represent write and read operations, respectively; dotted arrows represent a sync or async request-response communication between components. Components are colored according to the data they produce: raw data  , ML-specific data  , source code  , executable artifacts  , ML models  , ML training metadata  , ML model predictions  , and ML model metrics  . Remaining standalone components  orchestrate the execution of others. Components are also grouped into subsystems, with their background colored according to the students responsible for their development. Numbers  , roman numerals  and letters  are used as labels throughout section 2.
2.1 Data Collection
Creating an ML model that can detect respiratory insufficiency via voice is a supervised machine learning classification problem [2]. As such, it requires labeled data. The SPIRA team supports a Data Collection App (1) provided to volunteer data collectors to collect voices inside hospitals, where people with respiratory insufficiency may be found. The Data Collection App (1) sends data to a Data Collection API (2), which in turn stores audio in an Audio Key-Value Database (A), and saves patients’ info in a Document Database (B).
After each data collection, data collectors register a unique identifier for each participant. This ID can be cross-referenced with data shared by partner hospitals via 3rd-party Hospital APIs (3). This way, researchers can create a Label Store (C) that provides a ground truth for training models.
2.2 Continuous Training
Casanova et al.proposed a Deep Learning (DL) architecture based on Convolutional Neural Networks (CNNs) to detect respiratory insufficiency from audio signals [2]. As more data becomes available as ground truth, there is potential to retrain the model and improve its accuracy [9].
The proof-of-concept shared by Casanova et al. [2] can be divided into two components: a Feature Engineering Pipeline (I), to process audio signals for the training algorithm; and a Training Pipeline (II), to train and validate the CNN.
The development of these two pipelines is an opportunity to introduce standard machine learning design patterns [10, 3], such as a Feature Store (D), a Metadata Store (E), and a Model Registry (I). For the SPIRA project, the goal is to use MLFlow to make the two latter roles. On the other hand, for the features, the proposal is to use a simple key-value database like MinIO.
2.3 Development
As more data becomes available as ground truth, there is potential to redesign the architecture of the model used by SPIRA to improve its accuracy [9]. To make experimentation easy, it is desirable to provide a standardized environment for data scientists, compatible with production environments [16].
Currently, all SPIRA source code is open source in a Code Repository (J) at GitHub. Following the principles of Infrastructure as Code (IaC) [13], the SPIRA organization provides standard configurations for popular development tools used by data scientists and machine learning engineers. In particular, this includes a basic setup for Code Editors (8) (such as VSCode) or IDEs (8) (such as PyCharm), but can also include other utility scripts helpful for the developers.
2.4 Continuous Delivery
To support the creation of new versions of the MLES – after retraining or redesigning a model – automation is essential to make deployment seamless.
Taking advantage of GitHub, a Continuous Integration Pipeline (III) uses GitHub Actions to generate Docker containers for components, which are then deployed into the GitHub Container Registry, the de facto Artifact Store (K) for SPIRA. Conversely, the Continuous Delivery Pipeline (IV) and Continuous Delivery for Machine Learning Pipeline (V) use infrastructure configuration to deploy and rollback services in a Kubernetes-managed cluster.
2.5 Serving
Similarly to the Data Collection App (1), the SPIRA Client App (5) is meant to be used inside hospitals. However, it has a different goal: to provide (indirect) access to the SPIRA model, pre-diagnosing respiratory insufficiency via speech analysis. Therefore, its data collection is more critical.
Hospitals can have poor, unreliable internet reception. Multiple audios can be sent in a single burst whenever a user gets internet access. To prevent data loss, the Client App (5) must store all data collected locally before sending it to its corresponding Client API (6). Moreover, to prevent data inconsistency, the Client API (6) must validate all data received with the corresponding Client App (5) before storing it in its Audio Key-Value Database (G).
Deep Learning models such as the one proposed by Casanova et al.often require accelerator hardware (GPUs) to run efficiently [9]. Since executing the model is a costly operation, handling arbitrarily bursts of request is difficult.
To avoid this bottleneck, the SPIRA ML model can be executed in a separate Streaming Prediction Service (7), decoupled from the Client API (6). A Message Queue (H) stores prediction requests and corresponding results. Thanks to this separation, the Streaming Prediction Service (7) may be run in machines with GPUs, whereas other components are deployed in cheaper hardware.
2.6 Monitoring
Once the SPIRA MLES reaches production, it needs to be maintained in operation. The Telemetry Store (F) stores useful logs and data about its usage, while the Governance Application (4) summarizes statistics of its working status based on them. Combined, these two components help the SPIRA team to decide to retrain the ML model, or consider redesigning the ML model, or update and redeploy a component by triggering the Scheduler Service (9).
Figure 2: Incremental Development of the SPIRA Continuous Training Subsystem. The first version was published by Casanova et al. [2]. It has two main scripts: one to create datasets, and one to train the CNN model. Its code has an unplanned, organically developed structure, as it often occurs in experimental code developed by data scientists  [10, 16]. Therefore, the Experimental Architecture (v1) can be classified as a Big Ball of Mud [6]. The second version was created in the bachelor thesis of Daniel Lawand. It applies multiple design patterns [7] such as Chain of Responsability, Strategy, and Template Method to decouple different responsibilities within the pipeline. Moreover, the application logic was decoupled from external dependencies with the Ports and Adapters pattern [11]. Therefore, the Modularized Architecture (v2) can be classified as a Modular Monolith [15]. The third version was created in the bachelor thesis of Lucas Quaresma and Roberto Bolgheroni. It further expands on the second version by splitting the single application into multiple to facilitate the resilient execution of the pipeline in association with a Scheduler Service (9). Therefore, the Productionized Architecture (v3) follows a Microservices architectural style [14].
3 Incremental Implementation
The SPIRA MLES described in section 2 has been incrementally developed since 2021 via scientific initiation projects and bachelor theses supervised by M.Sc. Renato Cordeiro Ferreira and Prof. Dr. Alfredo Goldman at the University of São Paulo (USP). The scope of these projects is illustrated in figure 1.
In 2022, our team published the paper “SPIRA: Building an Intelligent System for Respiratory Insufficiency Detection” [4]. That experience report summarized challenges and lessons learned of developing the Data Collection and Serving subsystems, with contributions of the students Francisco Wernke and Vitor Tamae. This paper focuses on the Continuous Training subsystem, whose development process is summarized in figure 2.
4 Experimental Architecture (v1)
The first version of the Continuous Training pipeline was created as a proof-of-concept to showcase the SPIRA model. The code was published as part of the reproduction package [^2] https://github.com/Edresson/SPIRA-ACL2021 by Casanova et al.in 2020 [2].
The package has two main sets of scripts: one to create datasets (available in the scripts folder), and one to train the CNN model (available in the models folder). The code presents an unplanned, organically developed structure, as it often occurs in the experimentation phase by data scientists [10, 16]. Figure 3 illustrates an example of the Misplaced Responsability bad smell in the code, evidence of high coupling and low cohesion in the code. Therefore, this architecture can be classified as a Big Ball of Mud [6].
[…]
[ numbers=both, numbersep=0.5em, fontsize=, frame=lines, framesep=0.5em, highlightcolor=Violet!10, labelposition=topline, highlightlines= 2-7, 8-16, 18-20, 24-36, 32 , ]pythoncode/misplaced_responsability.py
Figure 3: Example of the Misplaced Responsibility bad smell. This code snippet from the Experimental Architecture (v1) shows part of the constructor of the auxiliary class Dataset. Lines 2-7 handle random number generation. Lines 8-16 assign values to attributes. Lines 18-20 and 24-27 handle assertions. Line 32 handles data loading from a CSV. The high coupling and low cohesion makes it difficult to create new experiments.
5 Modularized Architecture (v2)
The second version of the Continuous Training pipeline was created during the bachelor thesis research of Daniel Lawand, co-author of this paper. The v2 code is available open source at GitHub [^3] https://github.com/spirabr/SPIRA-training-prototype.
The system applies multiple design patterns [7] to increase the flexibility of the code. Some examples include: Chain of Responsability, to allow to dynamically structure the steps of the pipeline depending on the experiment that will be executed; Strategy, to allow choosing different data preprocessing, feature engineering, and model evaluation techniques; and Template Method, to allow the reuse of generic code snippets that can be extended for future experiments (often in combination with the Strategy pattern).
In another level of abstraction, the business logic was decoupled from external dependencies by using the Ports and Adapters architectural pattern [11], using Dependency Injection to connect different layers of the application. Therefore, the v2 architecture can be classified as a Modular Monolith [15].
6 Productionized Architecture (v3)
The third version of the Continuous Training pipeline was created during the bachelor thesis research of Lucas Quaresma and Roberto Bolgheroni, co-authors of this paper. The v3 code is available open source at GitHub [^4] https://github.com/spirabr/SPIRA-training.
The system expands on the Modularized Architecture (v2) by splitting the single application into multiple applications to facilitate the resilient execution of the pipeline in association with a Scheduler Service (9). If a failure occurs in one stage of the pipeline, it does not require re-executing it from start. Therefore, the v3 architecture follows a Microservices architectural style [14].
Testability was a key priority in this version. Following testing practices for microservices [15], the business logic received unit test, developed using Test-Driven Development (TDD), while external dependencies received integration tests, developed around the Ports and Adapters [11].
7 Challenges
The creation of the Experimental Architecture (v1) followed a typical data science workflow. Their goal is to create a proof of concept that demonstrates a problem can be solved using ML techniques [12]. In these workflows, serving a model, i.e., going into production, happens only on the last step of the process.
The Machine Learning Engineer (MLE) emerged as a specialization of a software engineer whose responsibility if to help to productionize such a model, helping the data scientist to think about the robustness of their model, and how it will fit an ML-Enabled System as a whole [1, 16].
Within this context, there were two main challenges during the evolution of the SPIRA Continuous Training architecture.
Separation of Concerns.
The general lack of cohesion in the v1 architecture made it extremely difficult to reuse its code. The process to create the v2 architecture became an exercise of Software Archeology [8], in which the code had to be read line by line to decipher intentions, since there were no clear abstractions and some responsibilities were spread throughout the code.
Testability.
Even after the introduction of patterns in the v2 architecture, in particular the Ports and Adapters pattern, testing was still challenging. With many intents represented in the original v1 architecture, it was difficult to recover what behavior should be tested for long-term maintainability once the v3 architecture was implemented.
8 Lessons Learned
The development of ML systems is inherently difficult, and developing them without planning makes their complexity even greater. There were two main lessons learned during the process of migrating the architecture.
Collaboration of Data Scientists and ML Engineers.
By bringing these two roles together since the beginning of the experimentation process, it is easier to design a more structured code that can be maintained long term.
Testability as a first-class concern.
If the code is planned with tests and validation since its inception, it is easier to create new experiments in different components of the pipeline.
9 Conclusion
This paper reported the experience of incrementally developing the Continuous Training subsystem of SPIRA, a project that aims to develop a MLES to detect respiratory insufficiency via speech analysis. Its architecture evolved in three stages: from a proof of concept Big Ball of Mud to an organized Modular Monolith to a tested set of Microservices. By learning from this process, other projects may avoid similar pitfalls to achieve production-readiness sooner.